OpenStudy (anonymous):

Three more than 4 times a number is equal to 5 less than 6 times the number. What is the number?

OpenStudy (anonymous):

4n+3=6n-5

OpenStudy (anonymous):

solve for n

OpenStudy (anonymous):

i got that already

OpenStudy (anonymous):

Okay, as long as you get that part. The answer is 4. 4x4= 16. 16 + 3 = 19 6x4=24. 24 - 5 = 19

OpenStudy (gorv):

Three more than 4 times a number is equal to 5 less than 6 times the number. let be number =x 4 times the number =4x three more =4x+3=6*x-5 3+5=6x-4x 8=2x x=4

OpenStudy (anonymous):

Yes. I just input the answer to show you it was correct.